At year-end 2016, Wallace Landscaping’s total assets were $1.6 million, and its accounts payable were $445,000. Sales, which in 2016 were $2.2 million, are expected to increase by 25% in 2017. Total assets and accounts payable are proportional to sales, and that relationship will be maintained. Wallace typically uses no current liabilities other than accounts payable. Common stock amounted to $445,000 in 2016, and retained earnings were $340,000. Wallace has arranged to sell $115,000 of new common stock in 2017 to meet some of its financing needs. The remainder of its financing needs will be met by issuing new long-term debt at the end of 2017. (Because the debt is added at the end of the year, there will be no additional interest expense due to the new debt.) Its net profit margin on sales is 6%, and 35% of earnings will be paid out as dividends.

a. What was Wallace's total long-term debt in 2016? Do not round intermediate calculations. Round your answer to the nearest dollar.

b. What were Wallace's total liabilities in 2016? Do not round intermediate calculations. Round your answer to the nearest dollar.

c. How much new long-term debt financing will be needed in 2017? (Hint: AFN - New stock = New long-term debt.) Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

a). Liabilities = Assets - Equity = 1,600,000 - 445,000 - 340,000 = $815,000

Long-term debt = Liabilities - Current liabilities

= $815,000 - $445,000 = $370,000

b). Total Liabilities = Long-term Debt + Accounts Payable

= $370,000 + $445,000 = $815,000

c). Additional Funds Needed = [A0 x (ΔS / S0)] - [L0 x (ΔS / S0)] - [S1 x PM x b]

Where,
Ao = current level of assets
Lo = current level of liabilities
ΔS/So = percentage increase in sales i.e. change in sales divided by current sales
S1 = new level of sales
PM = profit margin
b = retention rate = 1 - payout rate

AFN = [$1,600,000 x 0.25] - [$445,000 x 0.25] - [($2,200,000 * 1.25) x 0.06 x (1 - 0.35)]

= $400,000 - $111,250 - $107,250 = $181,500

New Long-term Debt = AFN - New Common Stock = $181,500 - $115,000 = $66,500